OpenStudy (anonymous):

Part 1: Solve the following quadratic equation using one of the methods listed below. Part 2: Using complete sentences, explain why you chose the method you used. Equation: 0 = x2 + 6x + 3 Methods: Completing the Square Factoring Quadratic Formula Graphing

OpenStudy (blockcolder):

I'd complete the square on this one, the reason why becomes apparent: \[x^2+6x+3+6=6\\ x^2+6x+9=6\\ (x+3)^2=6\\ x+3=\pm\sqrt6\\ x=-3\pm\sqrt{6}\] x^2+6x+3 is very close to a perfect square (x^2+6x+9), so I chose to complete the square.

OpenStudy (anonymous):

just cuz i suck at completing the square i'm gonna use the quadratic formula

OpenStudy (superhelp101):

lol this was posted 2 years ago

OpenStudy (anonymous):

hahaha. thats funny not really but ok. didn't see that. U just madde me crack up like crazy!

OpenStudy (superhelp101):

But please don't feel bad. You helped me out, the important thing is that we worked it out together.

OpenStudy (anonymous):

yep:)